# Error correction of parity-encoding-based annealing through post-readout decoding

### 13 Feb 2024 v1

@@ III.1 Classical repetition code
if we neglect the imbalance for the reliability of the estimator $r_{i}$. Equation (73) implies that the one-step MVD can correct up to $\left\lfloor\frac{N-1}{2}\right\rfloor$ errors. Because a set $\boldsymbol{A}$ of $J=d_{min}-1$ syndromes orthogonal on error $e_{1}$ of the logical spin state can be formed, one-step MVD can achieve the error-correcting capability $t_{RP}$ of the repetition code. Therefore, the repetition code can be completely orthogonalized in one step Massey (1963, 1968); Laferriere (1977); Clerk and Cain (1981); Shu and Costello (2004).
@@ V Conclusions
Subsequently, we addressed two classical error-correction schemes used for QA, namely the QAC and PE schemes, and investigated their derivation from the above concept. In QAC, the classical repetition code is merged to the logical Hamiltonian and encoded into the Hamiltonian describing a larger system composed of many copied systems (replicas). These replicas are mutually independent if the penalty Hamiltonian is omitted; otherwise, they are mutually correlated. It is obvious that if the copies are independent, the QAC executes sampling and measurement for independent replicas in parallel in a single-shot operation, instead of repeating the same operation on a single copy. The results should be consistent with the measurements of repeated experiments using a single copy. If the penalty term is introduced, it should produce a positive correlation among the replicas, which may be utilized to improve the sampling efficiency. This is very similar to extended ensemble algorithms, such as replica exchange and parallel tempering, in classical statistical physics IBA (2001) as well as simulated QA using the path-integral Monte Carlo method Marto?ak et al. (2002); Heim et al. (2015); Waidyasooriya and Hariyama (2020); Hu and Wang (2021), where the dynamics of an ensemble comprising the mutually interacting replicas are considered. In the PE scheme, soft annealing provides the encoded physical Hamiltonian directly, but its interpretation is not necessarily straightforward compared with that of the QAC.
@@ Acknowledgements.
The author acknowledges Dr. Yuki Susa, Dr. Yuki Kobayashi, Dr. Ryouji Miyazaki, Dr. Akihiro Yatabe, and Dr. Masayuki Shirane of NEC Corporation for their helpful discussions and comments. This paper is partly based on the results obtained from a project, JPNP16007, commissioned by the New Energy and Industrial Technology Development Organization (NEDO), Japan.

### 28 Feb 2024 v3

@@ III.1 Classical repetition code
if we neglect the imbalance for the reliability of the estimator $r_{i}$. Equation (73) implies that the one-step MVD can correct up to $\left\lfloor\frac{N-1}{2}\right\rfloor$ errors. Because a set $\boldsymbol{A}$ of $J=d_{min}-1$ syndromes orthogonal on error $e_{1}$ of the logical spin state can be formed, one-step MVD can achieve the error-correcting capability $t_{RP}$ of the repetition code. Therefore, the repetition code can be completely orthogonalized in one step Massey (1962, 1968); Laferriere (1977); Clerk and Cain (1981); Shu and Costello (2004).
@@ V Conclusions
Subsequently, we addressed two classical error-correction schemes used for QA, namely the QAC and PE schemes, and investigated their derivation from the above concept. In QAC, the classical repetition code is merged to the logical Hamiltonian and encoded into the Hamiltonian describing a larger system composed of many copied systems (replicas). These replicas are mutually independent if the penalty Hamiltonian is omitted; otherwise, they are mutually correlated. It is obvious that if the copies are independent, the QAC executes sampling and measurement for independent replicas in parallel in a single-shot operation, instead of repeating the same operation on a single copy. The results should be consistent with the measurements of repeated experiments using a single copy. If the penalty term is introduced, it should produce a positive correlation among the replicas, which may be utilized to improve the sampling efficiency. This is very similar to extended ensemble algorithms, such as replica exchange and parallel tempering, in classical statistical physics Iba (2001a) as well as simulated QA using the path-integral Monte Carlo method Martoňák et al. (2002); Heim et al. (2015); Waidyasooriya and Hariyama (2020); Hu and Wang (2021), where the dynamics of an ensemble comprising the mutually interacting replicas are considered Iba (2001b); Campillo et al. (2009); Cappe et al. (2004). In the PE scheme, soft annealing provides the encoded physical Hamiltonian directly, but its interpretation is not necessarily straightforward compared with that of the QAC.
@@ Acknowledgements.
The author acknowledges Dr. Yuki Susa, Dr. Yuki Kobayashi, Dr. Ryouji Miyazaki, Dr. Akihiro Yatabe, Dr. Tomohiro Yamaji, and Dr. Masayuki Shirane of NEC Corporation for their helpful discussions and comments. This paper is partly based on the results obtained from a project, JPNP16007, commissioned by the New Energy and Industrial Technology Development Organization (NEDO), Japan.